OpenStudy (anonymous):

5^4x-1=5^x+2 show me how to solve it pleaseeeee

OpenStudy (anonymous):

\[5^{4x+1}=5^{x+2}\] that show's it better

myininaya (myininaya):

5^(4x)-5^x-1-2=0 5^(4x)-5^x-3=0 let u=5^x so u^4=5^(4x) so we have u^4-u-3=0 lets see if we can use synthetic division to find any rational zeros if there are any we need to look at plus or minus 1 and plus or minus 3 so lets try -1 -1| 1 0 0 -1 -3 | -1 1 -1 2 ------------ 1 -1 1 -2| -1 lets try 1 1 | 1 0 0 -1 -3 | 1 1 1 0 _______________ 1 1 1 0 | -3 so neither 1 or -1 works lets try 3 3| 1 0 0 -1 -3 | 3 9 27 78 _______________ 1 3 9 26| 75 3 doesn't work how about -3 -3| 1 0 0 -1 -3 | -3 9 -27 84 ____________________ -3 -3 9 -28| 81 -3 doesn't work so there are no rational solutions we can use newton's method do you know newton's method but this method will only give us approximate answers

myininaya (myininaya):

u=u0-f(u0)/f'(u0) you know this method?

myininaya (myininaya):

which problem are we doing the first or second problem you posted?

OpenStudy (anonymous):

One sec let me read all that I'm totally lost :o

OpenStudy (anonymous):

The answer to the 5^4x+1=5^x+2 = 1

myininaya (myininaya):

\[5^{4x}-1=5^{x}+2\] i was doing the first problem you posted

OpenStudy (anonymous):

answer is, x=1

myininaya (myininaya):

no there are no rational zeros

OpenStudy (anonymous):

Oh okay I put that in wrong it's supposed to be \[5^{4x+1}=5^{x+2}\]

OpenStudy (anonymous):

the answer to that is x=1 But how do you get that?

myininaya (myininaya):

4x+1=x+2 4x-x=2-1 3x=1 x=1/3

myininaya (myininaya):

it isn't 1 it is 1/3

OpenStudy (anonymous):

maybe the teacher put it in wrong than thank youu

OpenStudy (anonymous):

I have more if you don't mind helping

myininaya (myininaya):

\[5^{4x-1}=5^{x+2}\] ?

myininaya (myininaya):

4x-1=x+2 4x-x=2+1 3x=3 x=1

OpenStudy (anonymous):

thank youu makes sense.

OpenStudy (anonymous):

\[9^{x+1}=27^{x-1}\]

myininaya (myininaya):

we don't have same base on both sides but 9=3^2 and 27=3^3 so we have \[3^{2(x+1)}=3^{3(x-1)}\]

myininaya (myininaya):

since we have the same base we need the exponents to be the same so we have 2(x+1)=3(x-1)

myininaya (myininaya):

2x+2=3x-3 2x-3x=-3-2 -x=-5 x=5

OpenStudy (anonymous):

that is right! yay!

OpenStudy (anonymous):

how about this one i'm stuck i worked some of it out

OpenStudy (anonymous):

\[3x-4 / 4-3x + 6-x / x-6 = 0\] I know the domain is 6, 4/3 the answer is no solution but how do you get that what I did was get like denominators which is (x-6)(4-3x) basically after doing all that i ended up with \[6x ^{4}-44x ^{2}+48=0\]

OpenStudy (anonymous):

on top of (x-6)(4-3x)

myininaya (myininaya):

\[\frac{3x-4}{4-3x}+\frac{6-x}{x-6}=0\] is this right?

OpenStudy (anonymous):

yesssss

OpenStudy (anonymous):

so then \[3x ^{2}-18x-4x+24+24-18x-4x+3x ^{?}\]

OpenStudy (anonymous):

\[3x ^{2}\] *

myininaya (myininaya):

ok so we know we cannot have zero on the bottom so we don't want 4-3x=0 and we don't want x-6=0 so we don't want x=4/3 and we don't want x=6 \[\frac{3x-4}{(-1)(3x-4)}+\frac{(-1)(x-6)}{x-6}=0\] \[\frac{1}{(-1)(1)}+\frac{(-1)(1)}{1}=0\] -1+-1=0 -2=0 which is never true so there are no solutions i see the way you are going about it that is longer lol

OpenStudy (anonymous):

Ah I'll go back to that one I have this huge final tomorrow at 7am! & All this stuff will be on it :(

myininaya (myininaya):

did it not make sense what i have above? you can go about it your way, but i think the way i above is much shorter? here is what you were attempting on this attachement let me know if there are any questions

OpenStudy (anonymous):

Oh yeah that helps! That's what I was doing. thanksss :d

OpenStudy (anonymous):

(b+5)(b+10)=6 x= -11, x= -4 How do you do that? :O

myininaya (myininaya):

b^2+10b+5b+50=6 b^2+15b+50-6=0 b^2+15b+44=0 (b+4)(b+11)=0 b=-4 b=-11

OpenStudy (anonymous):

wait you wrote it like this \[b ^{2+10b}\].. ?

OpenStudy (anonymous):

(b+5)(b+10)=6^2 + (5+10)b+(5)(10) ?

myininaya (myininaya):

i multiplyied b+5 and b+10

OpenStudy (anonymous):

like foil?

myininaya (myininaya):

yes